Suppose you sell a fixed asset for $128,000 when its book value is $148,000. If your company’s marginal tax rate is 21 percent, what will be the effect on cash flows of this sale (i.e., what will be the after-tax cash flow of this sale)?

ATCF
Answer #1

Loss on sale=148,000-128,000=$20,000

Hence aftertax cash flow=Sale proceeds+(Loss on sale*Tax rate)

=128,000+(20,000*21%)

=$132200